OpenStudy (anonymous):

Can you explain the equality of mixed partial derivites? I can take partial derivitives, but what does show that Fxy = Fyx??

OpenStudy (anonymous):

Verify that Fxy = Fyx F(x, y) = cos(xy) Fx = -ysin(xy) Fy = -xsin(xy)

ganeshie8 (ganeshie8):

\(\large F_{xy} = (-y\sin(xy))_y \) use product rule

ganeshie8 (ganeshie8):

\(F_{xy} = -[\sin(xy) + xy\cos(xy)]\)

ganeshie8 (ganeshie8):

\(F_{yx} = (-x\sin(xy))_x\) take the partial with.respect.to x

OpenStudy (kainui):

So basically all you're doing is you're taking a function, taking the derivative with respect to x first, then y second. Then start over with the same function, except this time take the derivative with respect to y first, then x. tadah, they're the same.

ganeshie8 (ganeshie8):

he may be looking for proof or something ? @jasonjohnson86

OpenStudy (usukidoll):

Fxy means take the partial derivative of y first and then take the partial derivative of x (if any)

OpenStudy (amistre64):

Fxy means (Fx)y ... partial wrt x first, then wrt y it is only when placed in some fractional format that it is read right to left:\[\frac{\delta F}{\delta x \delta y}\]would be y first then x .... cant recall how to code up the funny 'd'

OpenStudy (anonymous):

Thanks so much for the help everyone I got it!